If a test counts as 60 percent of my grade,how many points will be deducted from my average if I make a 42? What will my average be if I had a 69?
Answer by JulietG(1812)

If you already had a 69%, and that represents 40% of your grade, then you have 69 * 40 = 2760/10000
.
If the other 60% is 42, then that is 60 * 42, or 2520/10000.
Together you have 2520+2760 = 5280.
Move the decimal over to find that you now have 52.80%